OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue). A. 4/3 B. 2/3 C. 7/9 D. 4/27 I think the answer is A 4/3

OpenStudy (ledah):

seeing as it only says red and blue marbles I don't think it even cares about the other marbles. I always say go with your first answer. I think your right anyway.

OpenStudy (ledah):

but then again what IF it wants you to combine red and blue. then you got 7/9.

OpenStudy (ledah):

My guess is C because the red and blue total to 7 and the number of marbles in the bag is 9

OpenStudy (anonymous):

i'm not sure what they want

OpenStudy (anonymous):

I believe it's D because it's asking for 2 colors consecutively with replacement. \[\frac{\text{Number of Red Marbles}}{\text{Number of Marbles Total}} \times \frac{\text{Number of Blue Marbles}}{\text{Number of Marbles Total}}\]

OpenStudy (anonymous):

So, it's either C or D. Not completely sure.